Skip to article frontmatterSkip to article content
Site not loading correctly?

This may be due to an incorrect BASE_URL configuration. See the MyST Documentation for reference.

Sets, the sample space, and probability

Introduction

When we decide to study a topic, we often have in mind a population, a protocol to observe members of our population, a hypothesis that posits some aspects of our observed members. We collect data, we compile results. At the close of our study we want to know how the results ("what happened") contribute to our hypotheses ("what we thought may happen") and, importantly, the chances that if we reproduced the same study that the results would lead us towards the same conclusions.

For example, suppose we decide to study the impact of percutaneous intervention (PCI) versus optimal medical therapy (OMT) to treat patients how have had a myocardial infarction, often called a "heart attack". Patients are enrolled in a study and randomly assigned PCI or OMT. We hypothesize that PCI will result in a prolonged life from when the intervention took place onward. After our study we find patients who underwent PCI live on average 100 days longer than those who underwent OMT. However, it feels unsatisfactory to have a single number describing how effective PCI was at prolong life. We may want to know the chances PCI prolongs life 80 days, 30 days, or maybe zero days, compared to OMT.

We want to understand how likely we are to see specific events. The formal method of assigning chances to a set of events is called probability theory.

Probability lingo

Sets

We define a set as a collection of items, sometimes called elements. A set is typically given a capital letter (for example AA) and the elements are included inside curly braces.

Here, we define three sets

A={a,b,c}C={−1,1,3.14}Z={orange,1,puppydog}\begin{aligned} A &= \{ a,b,c\}\\ C &= \{ -1,1,3.14\}\\ Z &= \{ orange,1,puppy dog\} \end{aligned}

The elements inside a set are unordered and do not necessarily need to be numbers. A set can contain any object. The set AA could have been defined as {a,b,c}\{a,b,c\} or {c,b,a}\{c,b,a\}. We use the symbol ∈\in to express that an element "is a member of" a set. For example, a∈Aa \in A can be read "the element a is a member of the set A" or "the element a is in the set A". We can also communicate when an element is not a member of a set with the symbol ∉\notin. For example, c∉Cc \notin C is read "the element c is not a member of (not in) the set C".

Generating sets: We can define a set by enclosing in brackets each individual element. S={a,b,c,10}S = \{a,b,c,10\}. However, some sets may be easier to build with properties. A property is a function from elements to either true or false. For example we could define the property P(x)=1whenxisevenand0whenxisoddP(x) = 1 when x is even and 0 when x is odd. To build a set that contains elements with a specific property, we write S={x∣P(x)}S = \{x | P(x)\} or S={x∣xisaneveninteger}S = \{x | x is an even integer \}. This is read "the elements xx such that xx is an even integer". The vertical bar replaces the words "such that".

Two sets are equal is they contain the same elements. In other words, if for an element xx, x∈Ax \in A implies that x∈Bx \in B and if for an element yy, y∈By \in B implies that y∈Ay \in A then the set AA and set BB are equal. If a set AA and set BB are equal we write A=BA=B. If two sets do not contain the same elements then they are unequal and we write A≠BA \neq B.

Lets look at an example.

A={a,b,c}B={b,a,c}C={b,c}\begin{aligned} A &= \{ a,b,c\}\\ B &= \{ b,a,c\}\\ C &= \{ b,c\} \end{aligned}

Above, A=BA=B but A≠CA \neq C and B≠CB \neq C.

A set AA is a subset of a second set BB if all of the elements in AA are also elements in BB. We can say that AA is a subset of BB if every element in xx implies that element is in BB, or x∈Ax \in A implies x∈Bx \in B. We write A⊂BA \subset B to denote that AA is a subset of BB. To denote that a set AA is not a subset of BB, we write A⊄BA \not\subset B. In the example above, the CC is a subset of AA C⊂AC \subset A, but AA is not a subset of CC, or A⊄CA \not\subset C.

We can build new sets by operating on two or more existing sets.

Set intersection takes as input two or more sets, AA,BB, and returns a set that contains all the elements that are in both AA [and]{.underline} BB. We use the symbol "cap" to denote set intersection A∩BA \cap B and often say "A intersect B".

For the above sets AA and CC, their intersection is

A∩C={c,b}\begin{aligned} A \cap C = \{c,b\} \end{aligned}

because the element cc belongs to both set AA and set CC and because the element bb belongs to AA and CC.

We can intersect more than two sets.

Let A={a,b,10,12}A =\{ a,b,10,12 \}, Z={b,10,−1}Z =\{ b,10,-1 \}, U={a,b,c,d,10}U =\{ a,b,c,d,10 \}. Then

A∩Z={b,10}A \cap Z = \{ b,10 \}

and

A∩Z∩U={b,10}.A \cap Z \cap U = \{ b,10\}.

If AA is a subset of BB, A⊂BA \subset B, then the only elements both sets have in common are those in AA. The intersection of AA and BB is then

IfA⊂BthenA∩B=A.\begin{aligned} \begin{aligned} If A \subset B then\\ A \cap B = A. \end{aligned} \end{aligned}

Set union takes as input two or more sets and output a set that contains all the elements that belong to first set [or]{.underline} the second set [or]{.underline} the third set, and so on. We use the "cup" symbol to denote set union. As an example, consider the sets AA, ZZ, and UU to be the same as they are above. Then AA union ZZ is

A∪Z={a,b,10,12,−1}A \cup Z = \{a,b,10,12,-1\}

This is because each of the above elements belongs to at least one of the sets AA and ZZ. As another example,

A∪Z∪U={a,b,c,10,12,−1}A \cup Z \cup U = \{a,b,c,10,12,-1\}

Intersection and union are the most common set operations. Before we define the set operation complement, we need to discuss two special sets.

The universal set is the set of all elements. We denote the universal set with a G\mathcal{G}. The set which contains no elements is called the empty set, and we denote the empty set as ∅\emptyset.

If a set AA and set BB have no elements in common, then A∩B=∅A \cap B = \emptyset. We often say that the set AA and BB are disjoint. For example,

A={1,3,5}Q={2,4,6}A∩Q=∅\begin{aligned} \begin{aligned} A = \{ 1,3,5 \} \\ Q = \{ 2,4,6 \} \\ A \cap Q = \emptyset \end{aligned} \end{aligned}

Because the intersection between AA and QQ is empty, these sets are disjoint.

Now for our final set operation---complement. The set complement takes as input a single set, AA, and outputs a set with elements that are members of G\mathcal{G} but are not members of AA. We denote set complement in one of two ways: AcA^{c} or A′A'.

Let's look at an example of how these sets and set operations work. Define the universal set, $\mathcal{G}$, to be the set of all positive integers (i.e. $1,2,3,4,\cdots$). Lets define the sets $A_{1}=\{1\},A_{2}=\{1,2\},A_{3}=\{1,2,3\}, \cdots, A_{n}=\{1,2,3,4\cdots,n\}$. We use a subscript and a number to \"index\" our sets. An index is an easy way to keep track of sets (and many mathematical objects). $$\begin{aligned}
    \begin{aligned}
        A_{1} \cap A_{2} &= \{ 1 \}\\
        A_{1} \cap A_{5} &= \{ 1 \}\\
        A_{3} \cap A_{8} &= \{ 3 \}\\
    \end{aligned}
\end{aligned}$$ Note that $A_{3} \cap A_{8}$ is not the value $3$, but the set {3}. The examples above suggest a pattern for any set $A_{i}$ and $A_{j}$ where $i \leq j$: $$\begin{aligned}
    A_{i} \cap A_{j} = \{ i \}
\end{aligned}$$

Lets look at set union in this example $$\begin{aligned}
    \begin{aligned}
        A_{1} \cup A_{2} &= \{ 1,2 \}\\
        A_{1} \cup A_{5} &= \{ 1,2,3,4,5 \}\\
        A_{3} \cup A_{8} &= \{ 1,2,3,4,5,6,7,8 \}\\
    \end{aligned}
\end{aligned}$$ and we see the following pattern for $i \leq j$ $$\begin{aligned}
    A_{i} \cup A_{j} = \{1,2,3,4\cdots,j \} = A_{j}
\end{aligned}$$

Because we defined a universal set, we can look at set complement $$\begin{aligned}
    \begin{aligned}
        A_{1}^{c} &= \{2,3,4,5,\cdots \}\\
        A_{5}^{c} &= \{6,7,8,9\cdots \}\\
        A_{8}^{c} &= \{9,10,11,12,\cdots \}\\
    \end{aligned}
\end{aligned}$$

Applying set theory to probability

Foundation

The ideas about sets and set operations are the foundations for how we think about probability, about experiments, and hypotheses. We only need to recast the above set ideas to results from an experiment.

We use G\mathcal{G} to define the set of all possible outcomes from an experiment and call this set the sample space. The term "experiment" has a broad meaning. An experiment can mean everything from a randomized controlled trial to an observational study. Here an experiment is the process that generates outcomes.

An outcome is defined as an element of the sample space. An outcome is a single observation from an experiment, and we define an event as a set of outcomes. Most often we use oio_{i} to denote an outcome and EiE_{i} to denote an event.

The probability of an event EE is a mapping from the set EE to a number between, or equal to, the values 0 and 1. We use P(E)P(E) to denote the probability of event EE. We require that the probabilities assigned to individual sets consisting of a single element in EE add up to the probability of EE. Lets suppose there are nn outcomes in the event EE. Then

E={o1,o2,o3,⋯ ,on}P(E)=P({o1})+P({o2})⋯P({on})\begin{aligned} \begin{aligned} E &= \{ o_{1}, o_{2}, o_{3}, \cdots, o_{n}\} \\ P(E) &= P(\{o_{1}\}) + P(\{o_{2}\}) \cdots P(\{o_{n}\}) \end{aligned} \end{aligned}

We further require P(G)=1P(\mathcal{G}) = 1. In other words, the probability that something happens is certain. Note that we do not need to describe how we assign probabilities to events, we only describe what values we expect those numbers to be.

Lets further detail relationships between probabilities of sets that we would expect.

  1. If A⊂BA \subset B then P(A)≤P(B)P(A) \leq P(B) []{#subsetrule label=“subsetrule”}.

  2. P(A∪B)≤P(A)+P(B)P(A \cup B) \leq P(A) + P(B) (why?)

  3. P(A∪B)=P(A)+P(B)−P(A∩B)P(A \cup B) = P(A) + P(B) - P(A \cap B)

  4. If AA and BB are disjoint then P(A∪B)=P(A)+P(B)P(A \cup B) = P(A) + P(B)

The following three axioms (knowledge we assume true without proof) are called Kolmorogov’s axioms ()

  1. For any event EE, P(E)≥0P(E) \geq 0

  2. P(G)=1P(\mathcal{G}) = 1

  3. If E1E_{1} and E2E_{2} are disjoint then P(E1∪E2)=P(E1)+P(E2)P(E_{1} \cup E_{2}) = P(E_{1}) + P(E_{2})

Example: When we want to compute the probability that an outcome fall in any one of the events, say E1,E2,E3E_{1}, E_{2},E_{3}, knowing that these events have no outcomes in common, that they are disjoint, makes the computation easy. An intuitive way to think about events that are disjoint is if they cannot all occur at the same time. Suppose we wish to study the prevalence of stroke among patients who are older than sixety-five. Strokes can be categorized into transient ischemia attacks (TIA), ischemic stroke (IS), and hemorrhagic stroke (HS). Further lets assume the following probabilities for each event: P({TIA})=0.15P(\{TIA\}) = 0.15, P({IS})=0.75P(\{IS\}) = 0.75, P({HS})=0.10P(\{HS\}) = 0.10, and that these events cannot occur simultaneously. Then the probability of the event E={TIA,HS}E=\{TIA,HS\} can be broken into the union of two disjoint events E={TIA}∪{HS}E = \{TIA\} \cup \{HS\} and we can use what we know about these disjoint events to compute the probability of EE

P(E)=P({TIA,HS})=P({TIA}∪{HS})=P({TIA})+P({HS})      disjoint=0.15+0.75=0.90\begin{aligned} P(E) &= P(\{TIA,HS\} ) \\ &=P(\{TIA\} \cup \{HS\} ) \\ &=P(\{TIA\}) + P(\{HS\} ) \;\;\;disjoint \\ &= 0.15 + 0.75 = 0.90 \end{aligned}

The principle of equally likely outcomes---one way to assign probabilities

There are many different ways to assign probabilities to events. In this text, we will only consider using the principle of equally likely outcomes to assign probabilities. The principle of equally likely outcomes (PELO) states that the probability we assign to every event which contains a single outcome, Ei={oi}E_{i} = \{o_{i}\}, in a sample space G\mathcal{G} is equal.

We can use this principle to assign probabilities first to every individual outcome and then to arbitrary events. Assume the PELO is true, then

o1,o1,o1,⋯ ,on∈GEi={oi}E1∪E2∪E3∪⋯∪En=G    (why?)P(E1∪E2∪⋯∪En)=P(G)=1P(E1∪E2∪⋯∪En)=P(E1)+P(E2∪⋯∪En)    (disjoint)P(E1)+P(E2)+⋯P(En)=1n⋅p=1    (pistheconstantprob.weareafter)p=1/n\begin{aligned} o_{1},o_{1},o_{1}, \cdots,o_{n} \in \mathcal{G}\\ E_{i} = \{o_{i}\}\\ E_{1} \cup E_{2} \cup E_{3} \cup \cdots \cup E_{n} = \mathcal{G} \;\; (why?)\\ P(E_{1} \cup E_{2} \cup \cdots \cup E_{n}) = P(\mathcal{G}) = 1 \\ P(E_{1} \cup E_{2} \cup \cdots \cup E_{n}) = P(E_{1}) + P(E_{2} \cup \cdots \cup E_{n}) \;\; (disjoint)\\ P(E_{1}) + P(E_{2}) + \cdots P(E_{n}) = 1\\ n \cdot p = 1 \;\; (p is the constant prob. we are after)\\ p = 1/n \end{aligned}

The PELO tells us that the probability of an event with a single outcome is equal to one divided by the total number of outcomes in the sample space. The PELO also tells us that for a sample space with nn elements the probability of an event E={o1,o20,o4}E = \{o_{1},o_{20},o_{4}\} that contains three outcomes is P(E)=320P(E) = \frac{3}{20}.

PELO only works when we expect each outcome in our sample space to be equally likely.\
**Example 1** Suppose our experiment is a coin toss and we will observe whether or not the coin lands heads up (H) or tails up (T). If we assume the coin has not been altered in any way, then PELO applies. Our sample space is $\mathcal{G} = {H,T}$. Because there are two outcomes in our sample space, $P(\{H\}) = 1/2$ and $P(\{T\}) = 1/2$.

**Example 2** Suppose for our experiment we observe for one year a single patient who was admitted to the hospital because of an influenza infection and we plan to observe if that patient returns to the hospital because of a second infection. Our sample space contains one outcome $R$ if the patient is re-admitted to the hospital within one year and a second outcome if they are no re-admitted ($N$). Our sample space is $\mathcal{G} = {R,N}$. If we applied the PELO to this problem we would have to assume the probability this patient returns, or does not return, are equal. Intuitively,it feels unreasonable to assume these two events have equal probabilities.

We need an additional way to assign probabilities.

A frequentist approach to assigning probabilities

An empirical approach to assign probabilities to an event AA with sample space G\mathcal{G} is to do the following:

  1. Generate 1,2,3,4,⋯ ,N1, 2, 3, 4, \cdots, N outcomes from G\mathcal{G}

  2. Define a variable N(A)N(A) and set that variable to zero.

  3. If the 1st outcome is a member of AA than add one N(A)N(A), otherwise move on

  4. If the 2nd outcome is a member of AA than add one to N(A)N(A)
    ⋮

  5. If the Nth outcome is a member of AA than add one to N(A)N(A)

The above algorithm defined a variable N(A)N(A) that counts the number of outcomes that belong to AA out of the NN generated outcomes. We can assign to AA the probability

P(A)=N(A)NP(A) = \frac{N(A)}{N}

that is, the number of times we observed an outcome that belonged to A divided by the number of outcomes we observed.

As an example, suppose we want to understand the probability that patients with non-medically treated type II diabetes transition to needing treatment within one year of their diagnosis. Our outcomes are G\mathcal{G} = {"need treatment" and "no treatment"}. To assign a probability to the event E={needtreatment}E = \{need treatment\} using the frequentist approach we decide to request anonymized patient records from a set of hospitals, asking that each hospital only send those patients who had an original diagnosis of type II diabetes without medication and who have had a physician visit one year more later. After our data request, we received 5,000 patient records and find that 1,585 of these patients were asked to start medication for their type II diabetes within one year. A frequentist approach would assign

P(E)=1,585/5,000=0.317P(E) = 1,585/5,000 = 0.317

and also assign to the event F={notreament}F=\{no treament\}

P(F)=0.683    (why?)P(F) = 0.683 \; \; (why?)

.

Products, Conditionals, Baye’s Theorem, and Repeated Experiments

Product Sets

Suppose a novel vaccine was developed and we are asked to compare two probabilities: the probability a vaccinated patient is infected before 30 days after they receive their vaccine and the probability a patient who has not received a vaccine is infected before 30 days since they have enrolled in the experiment. To estimate these probabilities, we enroll 200 volunteer patients. We will assign 100 patients to be given the treatment and the remaining 100 patients will be observed without treatment.

Because of our experimental design, P({vaccinated})=1/2P(\{vaccinated\})=1/2 and P({notvaccinated})=1/2P(\{not vaccinated\}) = 1/2 with a sample space Gassignment\mathcal{G_{assignment}}, containing two outcomes {vaccinated,notvaccinated}\{vaccinated,not vaccinated\}. But our main interest is in the probability of infection within 30 days of enrolling in the study, and so our main interest is in the a related sample space Ginfection={infection,noinfection}\mathcal{G}_{infection} = \{infection, no infection\}. We need a way to combine these sample spaces together so that we can estimate the probability of a patient that is vaccinated [and]{.underline} that is infected.

Up until this point we have learned how to assign probabilities to events in one sample space. Lets look at how we may assign events to a combination of, and later a sequence of, sample spaces.

Products of sets: First, we’ll need to talk about the product of two sets. Let A={a,b,c}A = \{a,b,c\} and B={1,2,3}B = \{1,2,3\}. Then the Cartesian product C=A×BC = A \times B is a set where each element in AA is paired with each element in BB.

C={(a,1),(a,2),(a,3),(b,1),(b,2),(b,3),(c,1)(c,2)(c,3)}C = \{(a,1),(a,2),(a,3),(b,1),(b,2),(b,3),(c,1)(c,2)(c,3) \}

We use the notation (,)(,), called a "tuple", to denote pairs. A tuple is an outcome belonging to a sample space built from the Cartesian product of many individual sample spaces. Tuples are ordered. The tuple (a,1)≠(1,a)(a,1) \neq (1,a). The above is called a Cartesian product because you could imagine creating a grid where the horizontal axis has gridlines at "a", "b", "c" and the vertical axis has grid lines at "1", "2", and "3". Where the gridlines intersect are the tuples that belong to CC.

We can apply the Cartesian product to combine samples spaces. Define Gexperiment\mathcal{G}_{experiment} as the Cartesian product between Gassignment\mathcal{G}_{assignment} and Ginfection\mathcal{G}_{infection}, or

Gexperiment=Gassignment×Ginfection\mathcal{G}_{experiment} = \mathcal{G}_{assignment} \times \mathcal{G}_{infection}

This new sample space has the following outcomes:

Gexperiment={(vaccinated,infection),(vaccinated,notinfected),(notvaccinated,infection),(notvaccinated,notinfected)}\begin{aligned} \begin{aligned} \mathcal{G}_{experiment} = \{(vaccinated,infection),(vaccinated,not infected) \\ ,(not vaccinated,infection),(not vaccinated,not infected)\} \end{aligned} \end{aligned}

Our new sample space has 4 outcomes, two outcomes from the assignment sample space are paired with two outcomes from the infection sample space (2 outcomes ×\times 2 outcomes equals 4 outcomes in the new space).

With this new space in hand, we can assign probabilities to the outcomes "vaccinated [and]{.underline} infected" and the outcome "not vaccinated [and]{.underline} infected". You may hear events like the above called compound events, or joint events.

Let’s use the frequentist approach to estimate the probabilities of these four events. We enroll all 200 patients over a 6 months period and then follow each patient for 30 days. At 7 days, 14 days, 21 days, and at the end of there 30 day period they meet with a physician to relay information about how they feel that could indicate they had an infection. We collect the following data:

Vaccinated Infected Frequency Estimated Prob.


  Yes         Yes         20        20/200 = 0.1
  Yes          No         80        80/200 = 0.4
   No         Yes         40        40/200 = 0.2
   No          No         60        60/200 = 0.3

: Frequencies collected about our vaccination experiment

From our experimental evidence we would assign a probability of P({vacc,infected})=0.10P(\{vacc,infected\}) = 0.10 to those who were vaccinated and infected and a probability of P({novacc,infected})=0.20P(\{no vacc,infected\}) = 0.20 to those who were not vaccinated and were infected. But this was not what were were asked to compute. We were asked to compute the probability that someone is infected after they received ([given that they already had]{.underline}) a vaccine, and the probability of infection given a patient was not vaccinated. Probabilities of one event, given we have observed another event are called conditional probabilities.

Conditional probabilities

Assume a sample space G\mathcal{G}. We define the conditional probability of event AA given event BB as

P(A∣B)=P(A∩B)P(B)P(A|B) = \frac{P(A \cap B)}{P(B)}

Example: Lets use the definition above to compute the probability of an event AA given G\mathcal{G}, the sample space (remember that the sample space is a set and so technically an event). By the definition of conditional probbility

P(A∣G)=P(A∩G)P(G)P(A|\mathcal{G}) = \frac{P(A \cap \mathcal{G})}{P(\mathcal{G})}

We know that P(G)=1P(\mathcal{G}) = 1 (Kolmogorov axioms in [Kaxioms]{reference-type=“ref” reference=“Kaxioms”}) and so

P(A∣G)=P(A∩G)P(G)P(A∣G)=P(A∩G)1=P(A∩G)\begin{aligned} P(A|\mathcal{G}) &= \frac{P(A \cap \mathcal{G})}{P(\mathcal{G})} \\ P(A|\mathcal{G}) &= \frac{P(A \cap \mathcal{G})}{1}\\ &= P(A \cap \mathcal{G}) \end{aligned}

The set AA must be a subset of G\mathcal{G} and so by [subsetInter]{reference-type=“eqref” reference=“subsetInter”}

P(A∣G)=P(A∩G)=P(A)P(A|\mathcal{G}) = P(A \cap \mathcal{G}) = P(A)

The above example shows that we can think of the event that we condition on as a new sample space. Lets return to our example of vaccination and the incidence of infection.

To compute the probability of infection given a patient was vaccinated we need to compute

P(inf∣vacc)=P(inf∩vacc)P(vacc)P( inf | vacc ) = \frac{P(inf \cap vacc)}{ P(vacc)}

where the event "inf" contains those outcomes in Gexperiment\mathcal{G}_{experiment} which have "infected" in the fist position of their tuple (infected,⋅)(infected, \cdot) and the event "vacc" contains those outcomes which have "vaccinated" in the second position of their tuple (⋅,vaccinated)(\cdot, vaccinated ). Events which hold one or more positions in a tuple constant are often called marginal events and the associated probability is called a marignal probability. Then P(inf∩vacc)P(inf \cap vacc) are those outcomes with "infected" in the first position and "vaccinated" in the second position. We have a single outcome where this happens: {(infected, vaccinated)} and this outcome has a probability of P(inf∩vacc)=0.1P(inf \cap vacc) = 0.1.

Computing P(vacc)P(vacc) is only slightly more difficult. We can use what we learned about unions and disjoint events for help.

P(vacc)=P({(infected,vacc),(notinfected,vacc)})=P({(infected,vacc)}∪{(notinfected,vacc)})=P({(infected,vacc)})+P({(notinfected,vacc)})=0.1+0.4=0.5\begin{aligned} \begin{aligned} P(vacc) &= P( \{ (infected,vacc ), (not infected,vacc ) \} ) \\ &= P( \{ (infected,vacc )\} \cup \{(not infected,vacc ) \} ) \\ &= P( \{ (infected,vacc )\})+ P(\{(not infected,vacc ) \} )\\ &= 0.1+0.4 = 0.5 \end{aligned} \end{aligned}

We arrive at our final result.

P(inf∣vacc)=P(inf∩vacc)P(vacc)=0.10.5=0.2P( inf | vacc ) = \frac{P(inf \cap vacc)}{ P(vacc)} = \frac{0.1}{0.5} = 0.2

and we can do the same for

P(inf∣novacc)=P(inf∩novacc)P(novacc)=0.20.5=0.4P( inf | no vacc ) = \frac{P(inf \cap no vacc)}{ P( no vacc)} = \frac{0.2}{0.5} = 0.4

We report back to our experimental team that our estimated probability of infection from someone who received a vaccine is 0.2 and the probability of infection from someone who did not received a vaccine is 0.4, double the probability of a vaccinated individual.

Conditional probabilities give us, for free, an alternative way to compute the probability of the intersection of two sets. All we need to do is rearrange the definition of conditional probability

P(A∣B)=P(A∩B)P(B)P(A∩B)=P(B)P(A∣B)\begin{aligned} P(A|B) &= \frac{P(A \cap B)}{P(B)} \\ P(A \cap B) &= P(B)P(A|B) \end{aligned}

Equation [multrule]{reference-type=“eqref” reference=“multrule”} is called the multiplication rule. Lets explore how the multiplication rule may make computing probabilities easier.

Example: We want to compute the probability that it rains and we remember to leave the house with an umbrella. We can imagine our sample space having the following four outcomes: G\mathcal{G} = {(rain,remember),(rain,forgot), (no rain, remember), (no rain, forgot)}. We try to remember all the times that the outcome (rain,remember) occurred in our past---but we cannot. Instead, we can certainly estimate the probability we forgot our umbrella given that it rained because it was an unpleasant experience (for some). We estimate P(forgot∣rain)=0.95P(forgot|rain) = 0.95. To compute P(forgot∩rain)P(forgot \cap rain) all we need now is the probability it rains. We take a look at the weather station reports over the past year and use the frequentest approach to estimate the probability it rains:P(rain)=0.25P(rain) = 0.25. So then the probability it rains and we forget our umbrella is

P(forgot∩rain)=P(forgot∣rain)P(rain)=0.95×0.25=0.2375\begin{aligned} \begin{aligned} P(forgot \cap rain) &= P(forgot|rain)P(rain)\\ &=0.95 \times 0.25 = 0.2375 \end{aligned} \end{aligned}

The multiplication rule makes computations easier because we can think of events occurring in sequence. First it rains and second we forget our umbrella.

Conditional probabilities also allow us to express what it means for two events to be independent. An event AA is independent from an event BB if

P(A∣B)=P(A)P(A|B) = P(A)

We know the event BB occurred, but this has not changed the [probability]{.underline} that AA has occurred. When two events are independent we can compute their intersection more easily.

EventAandBareindependentP(A∩B)=P(A∣B)P(B)multiplicationruleP(A∩B)=P(A)P(B)independence\begin{aligned} \begin{aligned} &Event A and B are independent\\ &P(A \cap B) = P(A|B)P(B) &multiplication rule\\ &P(A \cap B) = P(A)P(B) &independence\\ \end{aligned} \end{aligned}

If two events are independent than the probability they occur together is the product of their probabilities. Statistical independence can be difficult to see.

Example We are recruited to track the evolution of an infectious agent for a team of public health officials (PHOs). To support future strategic planning the PHOs want to know the impact of intervention XX on increases or decreases in the incidence of this infectious agent. The PHO team collected for each county in their state whether the intervention was enacted, and whether the incidence of case counts of this infectious agent increased or decreased 60 days after the intervention was in place data.

Intervention Raise in infection Probability


   yes               yes              0.225
   yes                no              0.525
    no               yes              0.075
    no                no              0.175

: Probabilities estimated from our data using a frequentest approach.

When we look at this table, our first thought may be that when the intervention is enacted there is a high probability that the infection rate decreases (0.525), evidence that this is an important intervention for preventing the spread of this agent.

But be careful. Let’s compute the probability of intervention P(intervention)P(intervention) where we define the event intervention as {(intervention,rise),(intervention,norise)}\{ (intervention,rise), (intervention, no rise)\} and the probability of a raise in infection P(raise)P(raise) where the event raise is defined as {(intervention,rise),(nointervention,rise)}\{(intervention,rise),(no intervention, rise)\}. To do this we should define our sample space to be clear about what outcomes we can observe from our experiment.

Our sample space is the set of four pairs

G={(intervention,raise),(intervention,lower),(nointervention,raise),(nointervention,lower)}\begin{aligned} \begin{aligned} \mathcal{G}=\{ (intervention,raise),(intervention,lower)\\ ,(no intervention,raise) ,(no intervention,lower)\} \end{aligned} \end{aligned}

and we want to compute P(intervention)=0.225+0.525=0.75P(intervention) = 0.225 + 0.525 = 0.75 (why?) and P(raise)=0.225+0.075=0.30P(raise) = 0.225+0.075 = 0.30 (why?).

One way that can test whether intervention is independent of a rise or fall in incidence, is compare the probabilities we estimated from our data (in table 2{reference-type=“ref” reference=“tab.datafreq”}) to the product of individual events we computed above. Lets look at an example. The probability of an intervention and rise in incidence was estimated from our data to be P(intervention,raise)=0.225P(intervention,raise) = 0.225. Lets compare this probability to the product P(intervention)⋅P(raise)P(intervention) \cdot P(raise).

P(intervention)⋅P(raise)=0.75⋅0.30=0.225P(intervention) \cdot P(raise) = 0.75 \cdot 0.30 = 0.225

We have a match. Our estimated probability of an intervention and a rise occurring together is equal to the probability that an intervention occurs and a rise in incidence occurs. We can do the same procedure above for the remaining three scenario.

Intervention Raise Probability from data Probability assuming indep.


   yes         yes            0.225               0.75 (0.30) = 0.225
   yes         no             0.525               0.75 (0.70) = 0.525
    no         yes            0.075               0.25 (0.30) = 0.075
    no         no             0.175               0.25 (0.70) = 0.175

: Probabilities of the cooccurance of an intervention and the rise or fall of an infectious agent, and the probability of each of these four events assuming they are independent.

We see that the probabilities we collected form the data match the product of the probabilities of each individual event (Table 3{reference-type=“ref” reference=“tab.indep”}). Intervention is independent, does not change the probability, of a rise or fall in incidence of the infectious agent we are tracking. We will need to report these results to the PHO team and recommend they try alternative interventions to curb the spread of this agent.

The multiplication rule as a tool to compute sequential events

The multiplication rule equips us with a way to break down an event that involves many simultaneous phenomena into a sequence of, potentially easier to compute, probabilities.

Suppose we wish to study the impact diabetes and smoking may have on the probability of a stroke before the age of 50. We decide to observe a population of humans in the US who are all 18 years old and host annual check-ins to record whether each person has acquired diabates, is considered obese, and has experienced a stroke.

Our sample space is Ω={(x,y,z)∣  x∈{diab,  nodiab}andy∈{obese,  noobese}andz∈{stroke,  nostroke}}\Omega = \{ (x,y,z) |\;x \in \{diab, \;no diab\} and y \in \{obese,\; no obese\} and z \in \{stroke,\; no stroke\} \}. For example, one outcome in our sample is o=(diab,noobese,nostroke)o = (diab,no obese,no stroke) or the outcome that a patient has acquired diabates, is not considered obese, and has not experienced a stroke.

We may find it difficult to collect data that purposely recorded these three variables together, however it may be easier to find data that has paired diabetes with obesity and diabetes with stroke. The multiplication rule can give us some intuition about the outcomes we wish to study.

Define three events:

  1. D={(x,y,z)∣x=diab}D = \{ (x,y,z) | x=diab\}

  2. O={(x,y,z)∣y=obese}O = \{ (x,y,z) | y=obese\}

  3. S={(x,y,z)∣z=stroke}S = \{ (x,y,z) | z=stroke\}

We may find information to assign a probability to OO. Lets suppose we found P(O)=0.60P(O) = 0.60. Investigating further we find a dataset that compares those who have obesity and diabetes. This data may alow us to assign the following probabilities P(D∣O)=0.5P(D|O) = 0.5 and P(D∣Oc)=0.31P(D|O^{c}) = 0.31. Finally, we find a dataset that measured the prevalence of stroke among patients with/without diabetes and obesity, giving us the following probability assignments: P(S∣D∩O)=0.2P(S|D \cap O) = 0.2, P(S∣D∩Oc)=0.1P(S|D \cap O^{c}) = 0.1, P(S∣Dc∩O)=0.15P(S|D^{c} \cap O) = 0.15, and P(S∣Dc∩Oc)=0.01P(S|D^{c} \cap O^{c}) = 0.01.

If we feel that the above information applies to our study, we can compute the probability of simultaneous events by breaking them into a sequence of conditional probabilities. Suppose we wish to compute P(D∩O∩S)P(D \cap O \cap S). This can be separated using the multiplication rule:

P(D∩O∩S)=P(S∣O∩D)P(O∩D)=P(S∣O∩D)P(D∣O)P(O).\begin{aligned} P(D \cap O \cap S) &= P(S | O \cap D ) P(O \cap D)\\ &=P(S| O \cap D) P(D | O) P(O) . \end{aligned}

We can now compute the probability of our simultaneous event with the information we collected across several studies.

P(D∩O∩S)=P(S∣O∩D)P(D∣O)P(O)=0.2×0.5×0.60=0.06\begin{aligned} P(D \cap O \cap S) &= P(S| O \cap D) P(D | O) P(O)\\ &= 0.2 \times 0.5 \times 0.60 = 0.06 \end{aligned}

If we want, we can visualize all the conditional relationships related to DD, OO, and SS with a tree diagram. A tree diagram starts with a dot that represents the "root". We pick an event, say OO, and create two branches: one for when OO occurs and one for when OcO^{c} occurs.

We pick the next event, say DD, and create two branches from OO and two branches from OcO^{c} for a total of four branches. Conditional on OO, two branches represent the occurrence of DD and occurrence of DcD^{c}. Conditional on OcO^{c}, two branches represent the occurrence of DD and occurrence of DcD^{c}.

We can, in this example, do the same procedure for SS that we did for DD, but this time we need to look at the occurrence of SS and ScS^{c} for the four newest branches we created.

Partitions and the law of total probability

The multiplication rule is one way to use conditional probabilities to make computing probabilities of the intersection of many events more manageable. We can also use conditional probabilities to simplify computing single events easier.

A partition of the event AA is a collection of sets P={B1,B2,⋯ ,Bn}\mathcal{P} = \{B_{1}, B_{2}, \cdots, B_{n}\} such that A=(A∩B1)∪(A∩B2)∪(A∩B3)⋯∪(A∩Bn)A = \left(A \cap B_{1}\right) \cup \left(A \cap B_{2}\right) \cup \left(A \cap B_{3}\right) \cdots \cup \left(A \cap B_{n}\right) and for each i,ki,k pair of sets Bi∩Bk=∅B_{i} \cap B_{k} = \emptyset.

Figure 4:A partition 𝒫 = {B1, B2, B2} of the set A

Because BiB_{i} and BkB_{k} are disjoint for any choice of ii and kk, then A∩BiA \cap B_{i} and A∩BkA \cap B_{k} are also disjoint sets. We can use the disjoint property of our partition to break down the probability of AA into the sum of probabilities of AA intersect BkB_{k} for kk from 1 to nn.

P(A)=P((A∩B1)∪(A∩B2)∪(A∩B3)⋯∪(A∩Bn)=P(A∩B1)+P(A∩B2)++⋯+P(A∩Bn)(Disjoint)\begin{aligned} P(A) &= P(\left(A \cap B_{1}\right) \cup \left(A \cap B_{2}\right) \cup \left(A \cap B_{3}\right) \cdots \cup \left(A \cap B_{n}\right) \\ &= P\left(A \cap B_{1}\right) + P\left(A \cap B_{2}\right) + + \cdots + P\left(A \cap B_{n}\right) & (Disjoint) \end{aligned}

We can further manipulate the above using the multiplication rule to find

P(A)=P(A∩B1)+P(A∩B2)++⋯+P(A∩Bn)=P(B1)P(A∣B1)+P(B2)P(A∣B2)+⋯+P(Bn)P(A∣Bn)\begin{aligned} P(A) &= P\left(A \cap B_{1}\right) + P\left(A \cap B_{2}\right) + + \cdots + P\left(A \cap B_{n}\right) \\ &= P(B_{1}) P(A|B_{1}) + P(B_{2}) P(A|B_{2}) + \cdots + P(B_{n}) P(A|B_{n}) \end{aligned}

This is called the law of total probability. Intuitively the law of total probability says that the probability of an event AA can be computed by first considering the probabilities of a collection of related events, and then considering the probability that AA occurs given each related event.

Let G={(a,1),(a,2),(a,3),(b,4)}\mathcal{G} = \{(a,1), (a,2), (a,3), (b,4)\} and further suppose we assign the following probabilities to each outcome.

Outcome Probability


(a,1)(a,1) 0.10 (a,2)(a,2) 0.25 (a,3)(a,3) 0.14 (b,4)(b,4) 0.51

Define the events A={(a,1),(b,4),(a,2)}A = \{ (a,1), (b,4), (a,2) \}. The finest (as opposed to course) partition is a a collection of sets where each outcome in AA is in one, and only one, of the sets B1,B2,...B_{1},B_{2},.... For example a partition of AA could be P={B1,B2,B3}\mathcal{P} = \{ B_{1}, B_{2}, B_{3} \} where B1={(a,1)}B_{1} = \{(a,1)\}, B2={(b,4)}B_{2} = \{(b,4)\}, and B3={(a,2)}B_{3} = \{(a,2)\}. A courser partition is P={B1,B2}\mathcal{P} = \{ B_{1}, B_{2} \} where B1={(a,1)}B_{1} = \{(a,1)\} and B2={(a,2),(b,4)}B_{2} = \{(a,2),(b,4)\}.

Baye’s Theorem

We have investigated how the conditional probability is related to simultaneous events, gives us a natural definition of Independence, and can allow us to compute the probability of an event if we have a partition of that event.

The final relationship we will explore is between two conditional probabilities.

Baye’s Theorem relates two conditional probabilities to one another

P(A∣B)=P(B∣A)P(A)P(B)\begin{aligned} P(A | B) = \frac{P(B | A) P(A)}{P(B)} \end{aligned}

We can show that Baye’s theorem is true from the definition of conditional probability.

P(A∣B)=P(A∩B)P(B)definitionP(A∣B)=P(B∣A)P(A)P(B)multiplicationrule\begin{aligned} &P(A | B) = \frac{P(A \cap B)}{P(B)} & definition \\ &P(A | B) = \frac{P(B |A) P(A)}{P(B)} & multiplication rule \end{aligned}

Repeated experiments

Many natural experiments involve repeated observations. It may be of interest to observe the whether and assign a probability to "sunny" or "cloudy" weather conditions. However, many vacationers, who plan to spend 2, 3, or more days at a remote destination may want to know the probability that the next XX days are sunny.

The Cartesian product gives us a natural way to express repeated experiments. If we chose to repeat an experiment NN times where each experiment produced an outcome in G\mathcal{G}, we could imagine the set of outcomes as tuples of length NN where each entry in the tuple is from our sample space. In other words, a single outcome in this repeated experiment would be (o1,o2,o3,⋯ ,oN)(o_{1},o_{2},o_{3},\cdots,o_{N}) and this outcome is a member of the set G×G×G×⋯×G\mathcal{G} \times \mathcal{G} \times \mathcal{G} \times \cdots \times \mathcal{G}.

Back to our vacation example. If a vacationer wanted to understand the chances they had of enjoying 5 sunny days in a row, then we know that we could defined a sample space G={"sunny","notsunny"}\mathcal{G} = \{"sunny","not sunny"\} and we also know that the vacationer is interested in outcomes in the sample space

(d1,d2,d3,d4,d5)∈G×G×G×G×G(d_{1},d_{2},d_{3},d_{4},d_{5}) \in \mathcal{G} \times \mathcal{G}\times \mathcal{G}\times \mathcal{G}\times \mathcal{G}

where did_{i} is the outcome on day ii. We can frame the problem, and clearly layout the potential outcomes, for example on outcome is ("sunny", "sunny", "not sunny", "sunny", "not sunny").

The datapoint, dataset, and dataframe

The sample space, event, and outcome are all potential results from an experiment. When we conduct an experiment we will generate an outcome from our sample space and call this realized outcome a data point.

Consider the experiment of flipping a coin and recording whether the coin lands heads or tails side up. We can define a sample space Ω={H,T}\Omega = \{H,T\} where HH is an outcome that represents the coin landing heads up and TT represents tails up. Up until this point we have structured our experiment, but we have generated no data. We flip the coin and the coin lands tails side up. Now we have performed an experiment and generated the data point TT.

Now suppose that we conduct an experiment with the same sample space (Ω)(\Omega) a number NN times and with each experiment we record a data point. A tuple of data points dd is called a data set D=(d1,d2,d3,⋯ ,dN)\mathcal{D} = (d_{1}, d_{2}, d_{3}, \cdots, d_{N}) where did_{i} is the data point generated from the ithi^th experiment. We say that we have [drawn]{.underline} or that we have [sampled]{.underline} a data set D\mathcal{D}. Further, data points (d)(d) are often called [realized]{.underline} outcomes because they are no longer in a set of potential possibilities but are now determined items.

A data set D\mathcal{D} can be unwieldy depending on the number of data points, the complexity of the sample space, or both. A data frame is one way to organize a data set. A data frame F\mathcal{F} is a table where each data point dd in a dataset D\mathcal{D} is represented as a row in the table and if the data point is a tuple then a separate column is created for each position in the tuple.

Suppose we design an experiment to collect from humans predictions, two weeks ahead from the time of our experiment, of the number of incident cases and incident deaths at the US national level of COVID-19 (https://www.ncbi.nlm.nih.gov/pmc/articles/PMC7523166/). We decide to collect from each human whether they are an expert in the modeling of infectious disease, a prediction of incident cases, and a prediction of incident deaths. We draw a data set D\mathcal{D} of 50 human judgement predictions. We can organize this data set into a data frame

Expert Prediction of cases Prediction of deaths


Yes             145                    52
 No             215                    34
Yes             524                    48
Yes             265                    95
 No             354                    35

: Example data frame F\mathcal{F} built from a data set D\mathcal{D} that contains 5 data points where each data point is a tuple of length three.

Above, the first data point is (Yes,145,52)(Yes,145,52), the second data point is (No,215,34)(No, 215, 34), and so on until the last data point (No,354,35)(No, 354, 35). A data frame can also include informative information for others such as labels for each column.

Exercises

  1. $$\begin{aligned} A &= {1,2,3,4,5,6}; ;;B = {1,3,6} \ C &= {7}
    ;;D = \emptyset

    \end{aligned}$$

    1. Please compute A∩BA \cap B

    2. Please compute A∪CA \cup C

    3. Please compute A∪DA \cup D

    4. Please compute A∩DA \cap D

    5. Please compute (A∩B)∪(C∪D)(A \cap B) \cup (C \cup D)

  2. Let the sample space G={1,2,3,4,5,6,7}\mathcal{G} = \{ 1,2,3,4,5,6,7 \}

    1. Please compute AcA^{c}

    2. Please compute BcB^{c}

    3. Please compute DcD^{c}

    4. Please compute G∩A\mathcal{G} \cap A

    5. Is A⊂GA \subset \mathcal{G}?

    6. Is ∅⊂G\emptyset \subset \mathcal{G}?

  3. Delfina Szigethy---Class of 2025: 10 students in an Economics class and 10 students in a Biology class are asked to respond to the statement: Pick a number between 0 and 10. The data set collected for the economics class is E=(1,2,4,4,6,7,7,8,9,10)E = (1, 2, 4, 4, 6, 7, 7, 8, 9, 10) and the data set collected for the biology class is B=(0,2,2,5,6,6,7,9,9,10)B = (0, 2, 2, 5, 6, 6, 7, 9, 9, 10)

    1. Find the sample space Ω\Omega

    2. Form the set EE from the responses from the 10 students in economics

    3. Form the set BB from the responses from the 10 students in Biology

    4. Please compute E∪BE \cup B

    5. Please compute E∩BE \cap B

    6. Are the sets EE and BB disjoint?

    7. Suppose P(E)=0.3P(E) = 0.3. Please compute P(Ec)P(E^{c})

  4. Let the sample space Ω={0,1,2,a,b,c}\Omega = \{0,1,2,a,b,c\} and let A={0,1}A=\{0,1\}, B={x∣xisaletteroftheEnglishalphabet}B=\{x | x is a letter of the English alphabet\}

    1. Please compute A∩BA \cap B

    2. Please compute A∪BA \cup B

    3. Please compute AcA^{c}

    4. Is A∪B=ΩA \cup B = \Omega?

    5. If we assigned probabilities to all outcomes, could P(A∪B)=1P(A \cup B) = 1? why or why not?

  5. Let A=0,1,2A = {0,1,2} for some sample space G={0,1,2,3,4,5,6}\mathcal{G} = \{0,1,2,3,4,5,6\}. Further assume P(A)=0.2P(A) = 0.2.

    1. Are the sets AA and AcA^{c} disjoint? Why or why not.

    2. Simplify P(A∪Ac)P(A \cup A^{c}) into an expression that involves P(A)P(A) and P(Ac)P(A^{c}).

    3. Use Kolmogorov’s axioms to show that P(A)=1−P(Ac)P(A) = 1 - P(A^{c})

  6. Let Ω={x∣xisapositiveinteger}\Omega = \{x | x is a positive integer\}

    1. Are the sets ∅\emptyset and Ω\Omega disjoint?

    2. Simplify P(Ω∪∅)P(\Omega \cup \emptyset) into an expression that involves P(Ω)P(\Omega) and P(∅)P(\emptyset)

    3. Use Kolmogorov’s axioms to show that P(∅)=0P(\emptyset) = 0

  7. If A={1,2,3}A=\{1,2,3\} and B={2,3,4}B = \{2,3,4\} and C={1,3}C = \{1,3\}

    1. Can P(A)<P(B)P(A) < P(B)? Why or why not

    2. Can P(A)<P(C)P(A) < P(C)? Why or why not

  8. Use what you know about the intersection, about subsets, and about probability to show that P(A∩B)≤P(A)P(A \cap B) \leq P(A). Hint: How are A∩BA \cap B and AA related?

  9. Chinwe Okezie---Class of 2025: Let A={1,2,3,4},B={4,5,6}A = \{1,2,3,4\}, B=\{4,5,6\} and C={2,4,6}C = \{2,4,6\} and assume that the above outcomes follow the Principle of Equally Likely Outcomes.

    1. Can P(A)>P(B)P(A) > P(B)?

    2. Can P(B)>P(A)P(B) > P(A)?

    3. Can P(C)<P(B)P(C) < P(B)?

  10. Audrey Vitello---Class of 2025: Every year at the town fair there is a pie baking contest among the four best bakers. The best tasting pie wins and the outcome associated with this outcome is the baker’s number in the contest. For example, if baker one wins the outcome is the value 1, if baker two wins the outcome is the value 2 and so on.

    In the past ten years Mr. Brown, consistently contestant number 4, has won 3 times.

    1. Compute P(4)

  11. Suppose we wish to study the reemergence of cancer among patients in remition. We collect data on 1,000 patients who are in cancer remition and follow them for 5 years. At five years we are interested in the probability of a second cancer.

    1. Define a sample space G\mathcal{G} we can use to assign probabilities to a second cancer and no second cancer.

    2. After five years of followup we find that 238 patients experienced a second cancer. Use the frequentist approach to assign probabilities to a second cancer [and]{.underline} no second cancer.

    3. If you collected data on 2,000 patients do you expect the probability of a second cancer to change? How do you expect the probability to be different for 2,000 patients than with 1,000 patients?

  12. A study (link = https://www.science.org/doi/10.1126/science.abj8222 found that young adults were 32 times more at risk to develop multiple sclerosis (MS) after infection with the Epstein-Barr virus compared to young adults who were not infected by the virus. The experiment enrolled 10 million young adults and observed them for a period of 20 years.

    1. Design a sample space if we wish to study outcomes that describe the number of young adults who develop MS.

    2. Build the event (E1)(E_{1}) "less than 10% of young adults develop MS" using set notation.

    3. Build the event (E2)(E_{2}) "less than 5% of young adults develop MS" using set notation.

    4. Are E1E_{1} and E2E_{2} disjoint? Why or why not?

    5. Can P(E1)<P(E2)P(E_{1}) < P(E_{2})?

  13. Kareem Hargrove---Class of 2025: Suppose we wish to study the five year incidence of asthma among patients who were infected with SARS-CoV-2. We decide to enroll 10,000 patients who were infected with SARS-CoV-2, follow them for 5 years, and count the number of patients who were diagnosed with asthma at or before five years of followup.

    1. Define a sample space to describe whether a patient is or is not diagnosed with asthma within five years of followup.

    2. After the observation period ends we find 4,973 patients were diagnosed with asthma. Use the Frequentist approach to assign a probability to being diagnosed with asthma.

    3. Use the Frequentest approach to assign a probability to not being diagnosed with asthma.

  14. Please compute the following

    1. A={1,2,3}A = \{1,2,3\} and B={4,5,6}B = \{4,5,6\}. Please compute A×BA \times B (Answer should be a set of tuples)

    2. A={1,2,3}A = \{1,2,3\}. Please compute A×AA \times A (Answer should be a set of tuples)

    3. How many elements are in A×AA \times A (looking for a number)

    4. How many elements are in A×A×AA \times A \times A? (looking for a number)

    5. How many elements are in A×A×A×⋯×AA \times A \times A \times \cdots \times A where we take the Cartesian product NN times? (looking for a number)

  15. Define a sample space Ω={a,b,c,d,1,2,3,4,5}\Omega = \{ a,b,c,d,1,2,3,4,5 \} and let E1={1,3,5}E_{1} = \{1,3,5\}, E2={a,b,c}E_{2} = \{a,b,c\}, E3={a,d,5}E_{3} = \{ a,d,5 \}. We further assign the following probabilities

    Outcome P({Outcome})P(\{Outcome\})


      a            0.10
      b            0.05
      c            0.15
      d            0.02
      1            0.14
      2            0.25
      3            0.08
      4            0.04
      5            0.17
    1. Compute P(E1)P(E_{1}) (looking for a number)

    2. Compute P(E2)P(E_{2}) (looking for a number)

    3. Compute P(E3)P(E_{3}) (looking for a number)

    4. Compute P(E1∩E2)P(E_{1} \cap E_{2}) (looking for a number)

    5. Compute P(E1∩E3)P(E_{1} \cap E_{3}) (looking for a number)

    6. Compute P(E2∩E3)P(E_{2} \cap E_{3}) (looking for a number)

    7. Compute P(E1∣E2)P(E_{1} | E_{2}) (looking for a number)

    8. Compute P(E1∣E3)P(E_{1} | E_{3}) (looking for a number)

    9. Compute P(E2∣E3)P(E_{2} | E_{3}) (looking for a number)

    10. Compute P(E3∣E2)P(E_{3} | E_{2}) (looking for a number)

  16. Define a sample space Ω={(a,1),(b,1),(c,1),(a,2),(b,2),(c,2)}\Omega = \{ (a,1),(b,1),(c,1),(a,2),(b,2),(c,2) \} and let E1={(a,1),(a,2),(c,2)}E_{1} = \{(a,1),(a,2),(c,2)\}, E2={(c,2),(a,1)}E_{2} = \{(c,2),(a,1) \}, E3={(b,2)}E_{3} = \{(b,2) \}. We further assign the following probabilities

    Outcome P({Outcome})P(\{Outcome\})


    (a,1)          0.05
    (b,1)          0.22
    (c,1)          0.15
    (a,2)          0.02
    (b,2)          0.13
    (c,2)          0.43
    1. Compute P(E1)P(E_{1}) (looking for a number)

    2. Compute P(E2)P(E_{2}) (looking for a number)

    3. Compute P(E3)P(E_{3}) (looking for a number)

    4. Compute P(E1∩E2)P(E_{1} \cap E_{2}) (looking for a number)

    5. Compute P(E1∩E3)P(E_{1} \cap E_{3}) (looking for a number)

    6. Compute P(E2∩E3)P(E_{2} \cap E_{3}) (looking for a number)

    7. Compute P(E1∣E2)P(E_{1} | E_{2}) (looking for a number)

    8. Compute P(E1∣E3)P(E_{1} | E_{3}) (looking for a number)

    9. Compute P(E2∣E3)P(E_{2} | E_{3}) (looking for a number)

    10. Compute P(E3∣E2)P(E_{3} | E_{2}) (looking for a number)

  17. Imani Ashman---Class of 2024: A college student has been up studying for their midterm and decides to visit a coffee shop before the exam. The probability that they order a coffee (CC) is 0.2, the probability they order a bagel (B)(B) is 0.4, and the probability they order a bagel given that they have already ordered a coffee is 0.3. There is also the potential that they don’t order a bagel or a coffee before the midterm (N). The sample space is defined as G={B,C,N}\mathcal{G}= \{B,C,N\}.

    1. Compute P(B∩C)P(B \cap C)

    2. Compute P(C∣B)P(C | B)

    3. Compute the probability that they order nothing?

  18. Given two events AA and BB, show that P(A∣B)≥P(A∩B)P(A|B) \geq P(A \cap B) (looking for a short explanation and mathematical argument)

  19. Suppose we wish to study adverse outcomes among patients who have unprotected left main disease (Stone et al. (2019)). In this experiment (called a clinical trial) we randomize patients to receive percutaneous intervention (PCI) or a coronary artery bypass graft (CABG). We wish to study the number of patients who received a PCI [and]{.underline} experienced a myocardial infarction (MI) between the time they had their procedure and 5 years, however we only know three pieces of information: (i) the probability a patient was randomized to PCI was 0.5, (ii) the probability a patient was randomized to CABG was 0.5, and (iii) we know that the probability of a MI was 0.2 among patients who received a PCI.

    1. Define a sample space that will allow us to compute the probability a patient was randomized to PCI and experienced an MI (looking for a description of the sample space and a set)

    2. Please compute the probability a patient experiences an MI and was randomized to PCI. (looking for a number)

    3. Please compute the probability a patient does not experience an MI and was randomized to PCI. (looking for a number)

  20. If two events AA and BB are disjoint, are they also independent? (looking for a description and mathematical argument to backup your description.)

  21. Researchers studied epidemiological characteristics of a variant of SARS-CoV-2 called B.1.526 (Yang et al. (2022)). Suppose we too wanted to study the impact of B.1.526 on a population of patients who have been admitted to the hospital for COVID-19 by collecting each patient’s age, whether they have were vaccinated against COVID-19, and whether their infection was due to the B.1.526 variant.

    1. If we define our experiment as generating the above three pieces of information from a single patient, define a sample space (Ω\Omega) for these potential outcomes. (looking for a short description and the sample space)

    2. Define a sample space if we repeated our above experiment 2 times. In other words, what would be our sample space if we collected information from 2 patients?

  22. Can a data set ever have more elements than a sample space? Explain. (looking for a paragraph with some examples)

  23. Suppose we collect the following data about the co-occurence of patients admitted to the hospital for influenza-like illness (ILI) and whether the patient does or does not work in a clinical setting. After data collection we estimate the following probabilities: The probability that a patient works in a clinical setting is 0.12. The probability that a patient who works in a clinical setting is admitted to the hospital for ILI is 0.2, and the probability a patient who does not work in a clinical setting was admitted to the hospital for ILI is 0.3.

    1. Define a sample space to study outcomes related to the co-occurance of ILI/No ILI and patients who do/do not work in a clinical setting. (Looking for a set and a very brief description of an outcome or two.)

    2. Compute the probability a patient is admitted to the hospitals for ILI and they work in a clinical setting. (Looking for an equation and number)

    3. Compute the probability a patient is not admitted to the hospital for ILI and they work in a clinical setting. (Looking for an equation and number)

    4. Compute the probability a patient is admitted to the hospital for ILI and they do not work in a clinical setting. (Looking for an equation and number)

    5. Compute the probability a patient is not admitted to the hospital for ILI and they do not work in a clinical setting. (Looking for an equation and number)

  24. Matthew Jung---Class of 2025: Pretend you are on a game show. To win a prize, you are randomly assigned either door one, door two, or door three. If you are assigned door one, there’s a 9% chance to win $20,000, a 20% chance to win $10,000, and a 10% chance to win $40,000. If you are assigned door two, there’s a 10% chance to win $50,000 and a 30% chance to win $5,000. If you are assigned door three, there’s a 1% chance to win $100,000 and a 20% chance to win $15,000.

    1. Define the sample space for all of the possible outcomes.

    2. Build the table of all possible outcomes and their probabilities.

    3. Compute the probability of winning $40,000 given you get door two.

    4. Compute the probability of winning less than $40,000 given you get door one.

    5. Compute the probability of winning $100,000 given you get door three.

  25. Show that for two events AA and BB the P(A∣B)P(B)≤P(B)P(A|B)P(B) \leq P(B). Why is this the case intuitively? (Looking for two brief descriptions)

  26. Please use the above tree diagram to answer the following questions.

    1. Fill in the P(Oc)P(O^{c}) which corresponds to (A) in the figure

    2. Fill in the P(D∣Oc)P(D | O^{c}) which corresponds to (B) in the figure

    3. Fill in the P(Dc∣O)P(D^{c} | O) which corresponds to (C) in the figure

    4. Please compute P(D)P(D)

    5. Please compute P(Dc)P(D^{c})

  27. Walker Frisbie---Class of 2025: Suppose we want to study the relationship between the probability you feel ill after riding the roller coaster called the "Intimidator 305" (An actual roller coaster located in VA and my #1 favorite. I have ridden 150 different coasters) and the following three events: (i) we ate two hotdogs just before the ride, (ii) you rode this same coaster 10 times in the past 30 mins, and (iii) you have not had any water and are deydrated.
    Build a probability tree diagram with a root. Next add two branches that represent the events that you ride the Intimidator and that you do not ride the Intimidator. Given you ride the Intimidator with probability 0.9, create three branches for the events that you ate two hotdogs with probability 0.3, rode the coaster 10 times with probability 0.2, and are dehydrated with probability 0.5. Given you did not ride the Intimidator, create three branches for the events that you ate two hotdogs with probability 0.6, rode the coaster 10 times with probability 0.2, and are dehydrated with probability 0.2. For each of the six branches you created, add two branches: one for the event you feel ill and one for the event that you do not feel ill.

    1. Assign your own probabilities to each of these ill/not ill events.

    2. Compute P(feelill)P(feel ill)

  28. Influenza is a contagious virus that enters the respiratory system from a human’s nose, throat, or lungs. Common symptoms are cough, chills, and a fever. Suppose we collect data from a local hospital about those who were infected and not infected by the influenza virus and the above three symptoms. Let Ω={(x,y)∣xis"Flu"or"notFlu"andy∈{cough,chills,fever}}\Omega = \{ (x,y) | x is "Flu" or "not Flu" and y \in \{ cough ,chills , fever \} \}. Suppose further that we define events FF = { (x,y) | x is "Flu" }, CC = { (x,y) | y is "Cough" }, HH = { (x,y) | y is "Chills" } , EE = { (x,y) | y is "Fever" }, and that P(C∣F)=0.5P( C|F ) = 0.5 and P(C∣Fc)=0.15P( C|F^{c} ) = 0.15, P(H∣F)=0.25P( H|F ) = 0.25 and P(H∣Fc)=0.01P( H|F^{c} ) = 0.01, P(E∣F)=0.98P( E|F ) = 0.98 and P(E∣Fc)=0.35P( E|F^{c} ) = 0.35, P(F)=0.02P(F) = 0.02

    1. A patient presents with the chills. What is the probability they have influenza?

    2. What symptom from the above, if present, would indicate with a high probability a patient has influenza?

Student Advice, Examples, and Insights

Jessica Berman - Class of 2025-Advice: When I see a problem pertaining to conditional probability it helps me to talk the question out and actually say to myself “what is the probability of A given B has already occured” because P(A|B) is not clear to me without saying it aloud.

Jessica Berman - Class of 2025-Example: A fun fact about conditional probability is that it is used in politics. For example, if there are 4 candidates running in a presidential election, the chances of winning for each of them is 25

Monelli Esfandiary---Class of 2025---Insight: PELO assumes you assign the same probability to every outcome in your sample space. The frequentist approach assigns a probability to each outcome that is proportional to the number of times we observe that outcome in a dataset. Though the PELO and Frequentist approach appear different, we can show that the PELO is a special case of the Frequentist approach. Suppose we create a sample space G={H,T}\mathcal{G} = \{H,T\}. PELO would assign P({HH}) = CC and P({TT}) = CC, and so (1 = 2C2C) = (0.5 = CC). But we could arrive at the same result using the Frequentist approach if we sampled every outcome from G\mathcal{G} once.

Anyah Kumar---Class of 2025---Statistical Independence: The topic from chapter one that I liked the most was statistical independence. Two events, AA and BB, are statistically independent when one probability does not affect the probability of the other. This may remind you of mutually exclusive events. However, mutually exclusive events cannot be statically independent because when one event occurs the other event CANNOT occur. That’s why I like statistical independence, because the probability of one event doesn’t change the probability of another event which makes computing the probability of the two events simpler.

Ava Barrentine---Class of 2025---Baye’s thm: Baye’s Theorem is the most interesting to me as it shows the probability of an event, given another has already happened. I never thought one probability could be calculated based on a second event taking place. This theory is useful for so many reasons. For example, this equation can be used to calculate the probability of there being a forest fire when a gender reveal party has happened (probably very low). Say the probability of a forest fire (event AA) is P(A)=0.3P(A) = 0.3. The probability of a baby shower happening in/by the woods (BB) is P(B)=0.05P(B)= 0.05. The probability of a forest fire happening due to a baby shower P(A∣B)=0.007P(A|B) = 0.007. For Baye’s Theorem, we can compute : (0.3 x 0.007)/0.05 = 0.042 = 4.2% chance that given that a forest fire started, a baby shower took place near the woods.

*Josephine Osroagbo---Class of 2025---PELO:*I feel that the Principle of Equally Likely Outcomes does not occur frequently when defining an experiment and assigning probabilities to events. For example, suppose we wish to study the prevalence of COVID-19 hospitalizations among those who are 60 years of age or older with a vaccine to guard against infection and without a vaccine.

We can define a sample space for a single patient as G={(H,V),(H,Vc),(Hc,V),(Hc,Vc)}\mathcal{G} =\{(H,V), (H,V^{c}), (H^{c},V), (H^{c},V^{c})\} where HH represents the event that a person is hospitalized and VV represents that a person was vaccinated against infection by SARS-CoV-2.

If we applied PELO to this experiment we would have to assume the above four events occur with equal probability. Intuitively, this is likely not the case.

Neha Paras---Class of 2025---A connection between PELO and Uniform: While looking back at my chapter 1 notes, I grazed through the section about PELO. PELO states that the probability of each event in a sample space with a single outcome is equal. This reminded me of the class lesson where we were introduced to the uniform discrete distribution. I thought PELO and the uniform distribution were similar because they both assign to an event with a single outcome a probability of one divided by the number of outcomes in the sample space. I thought the connection between these two concepts was interesting, and it made me think more about how distributions fall outside of uniform. I am excited for future learning about the uniform distribution, for random events, as well as for the normal distribution.

Glossary

References
  1. Stone, G. W., Kappetein, A. P., Sabik, J. F., Pocock, S. J., Morice, M.-C., Puskas, J., Kandzari, D. E., Karmpaliotis, D., Brown, W. M., Lembo, N. J., Banning, A., Merkely, B., Horkay, F., Boonstra, P. W., van Boven, A. J., Ungi, I., Bogáts, G., Mansour, S., Noiseux, N., … Serruys, P. W. (2019). Five-Year Outcomes after PCI or CABG for Left Main Coronary Disease. New England Journal of Medicine, 381(19), 1820–1830. 10.1056/nejmoa1909406
  2. Yang, W., Greene, S. K., Peterson, E. R., Li, W., Mathes, R., Graf, L., Lall, R., Hughes, S., Wang, J., & Fine, A. (2022). Epidemiological characteristics of the B.1.526 SARS-CoV-2 variant. Science Advances, 8(4). 10.1126/sciadv.abm0300